RC Update and Run Plan for March 23, 2020

From HallCWiki
Jump to: navigation, search

Current RC: Jay Benesch benesch@jlab.org

RC Daily Meetings are M-F at 4:15 PM, 2nd floor Counting House Meeting Room Bluejeans Connection

General updates

  • The Accelerator is shutting down tomorrow (Tuesday), March 24 at 8am as we move to MedCon6
  • Follow the social distancing rules, use the disinfectant wipes to clean your keyboard and workspace for the upcoming shift crew.
  • JP is compiling a Target shutdown plan
  • We will have a final RC meeting Tuesday at 4:15pm

Current Run Plan

Run Plan for Beam: Monday SWING — Tuesday OWL

  • Take remaining Optics data as indicated in the table below
    • NOTE: The run time for the remaining setting assumes beam continues with the uptimes we've been seeing earlier today.
    • NOTE: The beam charge asymmetry does not matter for the optics runs!
    • Take and NMR every 2 hours to monitor spin up
  • Use maximum beam current that the Accelerator can provide and remain stable (up to a maximum of 45 uA on the Carbon target, so far 25-30 uA seems to be maximum)
  • Set prescales to get close to 3.5--4 kHz accepted trigger rate in each arm as needed
    • NOTE: If the un-prescaled trigger rates are low enough, then use 3/4 trigger instead of EL-REAL

Optics Kinematic Settings

P THETA P THETA (beam-on-target)
Optics: HMS KinB / SHMS KinY
Optics/No Sieve (Acceptance) 4 20 (-) 5.6 18 (-) C-Optics 40 4mm 0.5 0.05 ELREAL ELREAL DONE Need >100k good electrons
No Target/No Sieve 4 20 (-) 5.6 18 (-) No Target 40 4mm 0.5 0.05 ELREAL ELREAL SKIP Need >100k good electrons
Optics/Sieve (angle & ytar) 4 20 (-) 5.6 18 (-) C-Optics 40 0.5mm 5 0.11 ELREAL ELREAL Take 4--5 more hours Need >50k good electrons
No Target/Sieve 4 20 (-) 5.6 18 (-) No Target 40 0.5mm 1 0.11 ELREAL ELREAL SKIP Need >50k good electrons

Pol He3 Production

  • Go to No Target
  • Configure the spectrometers:
    • NOTE: You will need to go to MOL and then ramp back down to the these settings.
 HMS: theta = 13.5 deg, P = -4.2 GeV/c
SHMS: theta = 11.0 deg, P = -7.5 GeV/c
  • Update the target correction coil settings to match the above SHMS setting.
  • While the spectrometers are settling, ask MCC to go to tune beam and take some harp swipes on 3H07A and 3H07B.
    • Verify that the (unrastered) beam profile is a nominal 200--500 um in X and Y. Call Brad if you are unsure.
  • Take an NMR measurement (if you haven't done one recently)
  • Move to the Pol 3He Cell and ask MCC to set the raster for 4.0 mm
    • Use the current target polarization setting (180 deg)
  • Set trigger for 3/4 in both DAQs
  • Ask for 5 uA of beam and start runs in both arms
    • Trigger rates should be around 1 kHz or so (not too important for this step)
    • NOTE: Monitor the online GUI and verify that there is no scraping around the perimeter of the raster pattern in the Raster (Track Cut) histogram
  • IF things look good, then
    • Stop the runs, leave beam running
    • Set the prescales to use EL-Real (trigger 2) in both arms
    • Raise the current and adjust prescales to get close to 3.5--4 kHz accepted trigger rate in each arm
    • Take data on the Pol 3He target until Tuesday 8am
    • NOTE: The beam charge asymmetry does matter for these production runs! (Watch the stripcharts on the big monitor.)
  • Take an NMR measurement after beam turns off
  • Ramp the spectrometers down to zero current to help prepare the Hall for the shutdown.

Notes for the Future

  • At d2n startup, it would be a good idea to quantitatively check the dispersion at the target. I (D.Mack) don't think we need a tight spec, but if it's as sloppy as 1 cm/% (hence a decent spectrometer), then we're more at risk for things like how much the beam moves when the locks are off, or scraping from an off-energy tail on the beam. Jay:MQA3C13 can be adjusted to change horizontal dispersion.

Pending Issues

  • Target Ladder software must throw an error on a 'voltage' readback mismatch! [1]
    Get 3H07B Harp repaired [2] [3]; and a puzzle [4]
    https://logbooks.jlab.org/entry/3807201 Has Arun's note about the unfortunate reference cell alignment. He implies
    NOTE: "Shift Leader must request a beam that is shifted horizontally to the beam left side by 2 mm to prevent scraping."
    pNMR issues
    Software/driver issue needs to be resolved. Checkout during next non-production down time (Moller or SHMS fix)
    EPR problems 
    [5] saw some progress on Tuesday [6], but issues remain.
    Tentative plan now is to continue the d2n (and A1n) program by replacing the photo-diode that is placed directly on the pumping chamber (we have many spares).
    It would be good to write up a procedure to streamline this replacement procedure for us and RadCon
    Kepco power supply 
    PS commissioning/installation (Todd, Bill) (Bill got a filter design from the vendor and has ordered parts.)
    New filter (for 180 Hz) to install for current mode operation fix. Can always go to voltage mode if it doesn't work.
    NMR Calibration Constants 
    Verify that calibration constants and input values associated with NMR and other polarimetry are consistent and will be correct when the software is restarted
    [7] [8] [9] [10] [11] (Bill, Junhao, Arun)
    USB bridge 
    Contact with Target readbacks through AnywhereUSB USB bridge still drop out on occasion
    iochc10 issues 
    Green now: Rebooted 3/2/20 with Dave Gaskell assistance. OPs wished to reboot hung iochc10 [12], noted in December as well [13]
    This runs the Moller cryo controls and should not be rebooted without expert oversight (does not impact operations). D. Gaskell is looking into getting this addressed properly with OPs.
    Compton related items in Hall C beam steer-up procedure still cause confusion between MCC and Shift Crew

Opportunistic Access Jobs

None at the moment.

Standing Orders

 If beam goes away for more than 2h, or if during the downtime, MCC does operations
 that may change the beam tuning, then do [https://hallcweb.jlab.or/wiki/index.php/Beam_and_Raster_Checkout_for_A1n/d2n_(Short_Procedure)
 beam recovery procedure] (do YAG viewer part only if you are suspicious of
 the harp scan results)
  • NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell the Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)
  • Check the position locks are on.
  • Do your best to accommodate Hall A's request for a PITA scan promptly. It is OK to stop the present runs and start new ones in 5--10 minutes after they are done.
  • Check that CODA is running normally (had one strange hangup lately)
  • Check online replays shortly after starting runs, & full replay results too.
  • Make sure all the beam specs are followed (e.g. no raster size below 4mm on 3He cell).
  • Always take 120Hz of EDTM
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.

Bluejeans remote connection information

 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links